34. Use the following information for questions 31-34: The average GPA for all college students in the United States as of 2006 was 3.11.  We want to test to see if the GPA of students at Texas A&M is higher than the national average.  Suppose we survey 47 randomly selected students at Texas A&M and the average GPA is 3.27, with a standard deviation of 0.54.  Assuming all conditions are met, conduct a hypothesis test at the 0.01 significance level.

Which of the following below best describes the p-value?

a.

p value = 0.0424

b.

0.02 < p-value < 0.025

c.

p-value = 0.9788

d.

0.04 < p-value < 0.05

e.

p-value = 0.0212
Answer #1

Solution:

H0: = 3.11

H1: > 3.11

The test statistics t is given by ..

t =  

= (3.27 - 3.11)/(0.54/47)

= 2.031

Now

df = n - 1 = 47 - 1 = 46

One tailed right sided test

So ,

p value = 0.024

0.02 < p-value < 0.025

Option b. is correct